# Reconstructing Intelligible Speech from the Pressure Sensor Data in HVACs

### v1

@@ 7.6 Impact of Sound Volume
7.7 Impact of Speaker Distance
@@
We vary the distance of a speaker from 0 m (touching the pressure sensor’s input port) to 3 m from the target pressure sensor. The result is shown in Fig. 10 (Right) for LSD and NISQA-MOS for 500 Hz to 8 kHz upsampling for 60 dB audio. In acoustics, the energy of a sound wave radiating from a point source decreases as the distance increases following the inverse-proportional law [60]. Therefore, the improvement in LSD and NISQA-MOS decreases with increasing distance from the audio source to the pressure sensor. It is also clear from Fig. 10 (Right) that WaLi performs well up to 1 m distance. After 1 m, the reconstructed audio has severely degraded intelligibility.
7.8 Impact of Speaker Orientation
@@
We use a loudspeaker to simulate a speaker and place the loudspeaker at 0°(in front of the loudspeaker), 90°(right of the loudspeaker), and 180°(left of the loudspeaker) of the pressure sensor. The loudspeaker is 5 cm away from the pressure sensor and generates a sound of 60 dB. The results are shown in Fig. 11 for LSD and NISQA-MOS for 500 Hz to 8 kHz upsampling. The LSD and NISQA-MOS are higher at 0° orientation, as at 0°, the audio is directed to the pressure sensor ports with minimum loss, potentially creating a strong vibration. Performance is the same at both 90°and 180°orientations, but lower compared to 0°.
Figure 11: Impact of orientation of the pressure sensor on the reconstruction quality of audio from the pressure sensor data.
7.9 Fine-Tuning on Victim’s Ground Truth Audio
@@
Please note that WaLi can reconstruct the audio even if the speaker is different as WaLi is tested with different speakers, which are not seen by the model during the training time. This makes our attack model more flexible compared to prior works [10, 13], where prior works assume that clean ground-truth speech of the victim is available via a microphone or a phishing call. However, if the attacker could manage the victim’s {ground truth, pressure sensor data} pair, the accuracy of the models would be greatly improved by WaLi. To evaluate this, we incorporate a small amount of {ground truth, pressure sensor data} pair from the victim to fine-tune the pre-trained network. As shown in Table VII, the performance gradually improves if the amount of data from the victim increases. For example, with just one minute of data, we observe significant performance gains, and after five minutes of additional data, the performance gain is minimal. This proves that WaLi can be fine-tuned with the victim’s audio, if possible, in a stronger assumption of the attack model. Table VII is for 60 dB audio for 500 Hz to 8 kHz upsampling.
@@
TABLE VII: Evaluating WaLi after fine-tuning on victim speaker’s ground truth audio.
7.10 Testing With Less Than 500 Hz Sampling
@@
We test WaLi while reconstructing speech from 250 Hz sampling frequency of the pressure sensor data, shown in Fig. 12. As only a few pitch frequencies are present within 125 Hz bandwidth of the 250-Hz-sampled data, the reconstructed speech in Fig. 12(c) is not intelligible with a NISQA-MOS value of 0.48. Therefore, the attacker should have access to DPSs that have at least 500 Hz sampling for proper intelligibility. However, the 250-Hz-sampled data may be used for hot word detection (i.e., our future work).
@@
8 Limitation and Discussion
Feasibility of the Attack: Sections 7.6, 7.7, and 7.8 indicate that the distance, orientation, and volume of the target speaker can restrict the effectiveness of the eavesdropping attack using DPSs. For a successful attack, the target victim should be at least 1 m close to the pressure sensor, have a pressure level of 60 dB or more, and should be closer to 0°  orientation. This situation exists in numerous scenarios (see Section 4), where DPSs are often installed in room walls, near diffusers, or within ventilation grilles, where human occupancy is high. If a target victim speaker unknowingly comes close to the pressure ports and continues a private conversation, the attacker can easily eavesdrop on the victim’s conversation using our WaLi.
Sampling frequency limit: WaLi achieves intelligible speech recovery with unrestricted vocabulary when the sampling frequency of the pressure sensor is 500 Hz or higher. If the sampling frequency of the DPSs is less than 500 Hz, the bandwidth of the speech is lower than 250 Hz, which is extremely narrow for an intelligible speech reconstruction. Therefore, the reconstruction quality degrades significantly at a sampling frequency lower than 500 Hz. A possible direction is to find the inner relationship between the pitch and the low-order harmonics to make WaLi work for a sampling frequency lower than 500 Hz.
9 Countermeasures
Audio damping: The simplest method of preventing eavesdropping using DPSs is to damp the audio using a physical damping device at the input of the pressure ports. The damping device should be such that it does not affect the normal pressure measurement of the pressure sensors. However, it should dampen the audio significantly. One solution is to use a pressure sampling tube longer than 1 m to dampen the audio significantly without impacting the normal pressure measurement. Another solution is to enclose the pressure pickup device using a box-like enclosure filled with sound-damping foam to dampen the audio. Both of these countermeasures are cheap and easy to adopt in a critical infrastructure where confidentiality might be an issue.
Reducing sampling frequency of DPSs: Another intuitive solution is to reduce the sampling frequency of pressure sensors in HVAC systems. However, we need to keep in mind that often reducing sampling frequency is not possible in some specific applications, where small pressure changes must be detected accurately and quickly for a stable control process. Therefore, for these closely controlled applications, an audio damping device should be used instead of reducing the sampling frequency of pressure sensors.
10 Conclusion
In this paper, we expose a new speech threat that adversaries can recover intelligible audio up to 8 kHz from severely aliased pressure sensor data, having a sampling frequency greater than 500 Hz. If a target victim speaker unknowingly comes close to the pressure ports and continues a private conversation, the attacker can eavesdrop on the victim’s conversation using our WaLi. WaLi can reconstruct audio even if the speaker is different and are not seen by the model during the training time. This makes our attack model more flexible compared to previous work. Using our WaLi, an attacker can secretly listen natural conversation behind the wall which is least unexpected. Moreover, we comprehensively evaluate WaLi using five metrics that have not been done before for speech eavesdropping tasks.

### v2

@@ VI-E Phase Reconstruction in Noisy Conditions
VI-F Impact of Sound Pressure Level
We vary the SPL from 60 dB to 85 dB with a 5 dB increment. We inject the audio into the pressure sensor ports through the 1m long sampling tube from a 0.5 m distance. The result is shown in Fig. 10 (Left) for LSD and NISQA-MOS for 500 Hz to 8 kHz upsampling. We can see that the LSD and NISQA-MOS improve rapidly between 60 dB and 70 dB. However, after 70 dB, the increase in volume does not drastically change the low-resolution frequency components. Therefore, the improvement is not sharp after 70 dB; rather, a steady ascent occurs. The LSD is close to 1, and NISQA-MOS is close to 2 for a wide volume range from normal (i.e., 60 dB) to loud speech, indicating WaLi’s capability to reconstruct intelligible speech from normal to loud in real-world scenario.
VI-G Impact of Speaker Distance
@@
We vary the distance of a speaker from 0 m (touching the pressure sensor’s input port) to 3 m from the target pressure sensor. The result is shown in Fig. 10 (Right) for LSD and NISQA-MOS for 500 Hz to 8 kHz upsampling for 60 dB audio. The improvement in LSD and NISQA-MOS decreases with increasing distance from the audio source because of the inverse-proportional law [43]. It is also clear that WaLi performs well up to 1.2 m distance. After 1.2 m, the reconstructed audio has severely degraded intelligibility.
Figure 11: (Left, Middle) Impact of speaker orientation on the audio reconstruction quality. (Right) Impact of sampling tube length.
VI-H Impact of Sampling Tube Length
We vary the sampling tube length from 0 m (touching the pressure port) to 3 m with a 0.5 m increment for 60 dB audio and 500 Hz to 8 kHz upsampling. Fig. 11 (Right) shows that LSD and NISQA-MOS degrade with the increase of the tube length as sound energy attenuates with the increase of the tube length, similar to speaker distance evaluation in Section VI-G.
VI-I Impact of Speaker Orientation
@@
We place the speaker at 0°(in front), 90°(left), and 180°(right) of the pressure sensor at 0.5 m away and generates a sound of 60 dB. The results are shown in Fig. 11 (Middle) for LSD and NISQA-MOS for 500 Hz to 8 kHz upsampling. The LSD and NISQA-MOS are higher at 0° orientation, as at 0°, the audio is directed to the pressure sensor ports with minimum loss, potentially creating a strong vibration. Performance is the same at both 90°and 180°orientations, but lower compared to 0°with still intelligible quality (NISQA-MOS = 1.5).
TABLE VI: Multi-speaker environment for 500 Hz to 8 kHz upsampling.
[…]
$\sim$15dB $\sim$40dB $\sim$50dB $\sim$15dB $\sim$40dB $\sim$50dB
2 speakers 1.21 1.37 1.65 1.72 1.61 1.31
3 speakers 1.38 1.52 1.81 1.65 1.47 1.14
4 speakers 1.59 1.67 1.93 1.41 1.29 0.92
VI-J Impact of Multi-Speaker Environment
We separately collect another 300 minutes of pressure data from 20 volunteers (9 male and 11 female) with two (one target speaker and one interfering speaker), three (one target speaker and two interfering speakers), and four (one target speaker and three interfering speakers) multi-speaker setups. The target speaker has an SPL of $\sim$60 dB, and the interfering speakers have SPLs of $\sim$15, $\sim$40, and $\sim$50 dB at 0.5m distance. We retrain WaLi using these multi-speaker data pairs (see Table VI for results). We see that with the increase in interference speakers and their SPLs, the reconstruction performance degrades. The reason is that when interfering speakers have SPLs close to the target speaker, the problem transforms from speech reconstruction to a speaker separation problem. WaLi performs well for 15 dB multi-speaker setup but its performance degrades for >>15 dB multi-speaker setup as WaLi is not optimized for speaker separation tasks. One alternative approach is that we can use speaker separation algorithms [44] first and then use WaLi for speech reconstruction. The performance of this approach will largely depend on how well the speaker separation algorithm performs because we already show that WaLi works well for speech reconstruction.
VI-K Fine-Tuning on Victim’s Ground Truth Audio
@@
Please note that WaLi can reconstruct the audio even if the speaker is different as WaLi is tested with different speakers, which are not seen by the model during the training time. This makes our attack model more flexible compared to prior works [10, 13], where prior works assume that clean ground-truth speech of the victim is available via a microphone. However, if the attacker could manage the victim’s {ground truth, pressure sensor data} pair, the accuracy of the models would be greatly improved by WaLi. To evaluate this, we incorporate a small amount of {ground truth, pressure sensor data} pairs from the victim to fine-tune the pre-trained network. As shown in Table VII for 500 Hz to 8 kHz upsampling, the performance gradually improves if the amount of data from the victim increases. For example, with just one minute of data, we observe significant performance gains, and after five minutes of additional data, the performance gain is minimal. This proves that WaLi can be fine-tuned with the victim’s audio, if possible, under a stronger assumption of the attack model.
TABLE VII: Evaluation after fine-tuning on victim’s ground truth audio.
@@
VI-L Testing With Less Than 500 Hz Sampling
@@
We test WaLi while reconstructing speech from a 250 Hz sampling frequency of the pressure sensor data, shown in Fig. 14 in Appendix G. As only a few pitch frequencies are present within 125 Hz bandwidth of the 250-Hz-sampled data, the reconstructed speech in Fig. 14(c) is not intelligible with an NISQA-MOS value of 0.48. However, the 250-Hz or less than 250-Hz-sampled data may be used for hot word detection [45].
TABLE VIII: Ablation study for 500 Hz to 8 kHz reconstruction.
Row Model L $\downarrow$ ST $\uparrow$ P $\uparrow$ S $\uparrow$ N $\uparrow$
(i) Real-valued net 1.58 0.71 1.35 4.24 1.43
(ii) Remove CGAB 1.32 0.74 1.45 7.54 1.48
(iii) Transformer in bottleneck 1.34 0.75 1.41 7.77 1.65
CNN in bottleneck 1.43 0.73 1.38 7.94 1.69
(iv) Magnitude-only loss 1.33 0.73 1.32 7.44 1.59
WaLi (complex-valued + conformer+CGAB+complex loss) 1.24 0.75 1.52 8.78 1.78
VI-M Ablation study
To understand the different parts of WaLi, we do an ablation study (see Table VIII) in the following four avenues: (i) We compare complex-valued WaLi with the real-valued WaLi in noisy conditions. We see that complex-valued WaLi does better in phase reconstruction and outperforms real-valued networks in real-world noisy conditions (see Sections VI-D and VI-E for details). (ii) We remove CGAB from WaLi and observe that it does not perform well when CGAB is removed because of the lacking of capturing long-range correlations in time and frequency axes (see Section IV-E for details). (iii) We separately use a transformer and CNN at the bottleneck layer and observe that they do not perform well compared to conformer (refer to Section IV-B for the reasons). And (iv) We use magnitude-only losses and find that it does not perform well because magnitude-only losses can not reconstruct clean phases directly in noisy conditions.
VII Limitation and Discussion
Feasibility of the Attack: For a successful attack, the attack constraints are: distance $\leq$1.2 m, orientation near $0^{0}$, volume $\geq$60 dB, and sampling $\geq$500 Hz (see Sections VI-F VI-G, VI-L, and VI-I). This situation exists in numerous scenarios (see Sections III and V), where DPSs are often installed in corridors, near entrance and exit doors, near diffusers, or within ventilation grilles, in industrial facilities. We run an experiment in the cleanroom corridor (see Section V, Fig. 7, & 4) to calculate what percentage of real-world conversations meet all attack constraints. In our case, the DPSs are located on the corridor wall near the entrance doors and far away from the entrance doors. A volunteer counts for 3 days on how many people (they are unaware of this experiment) are doing normal talking ($\sim$60 dB) while passing through that corridor with $\sim\leq$1.2 m close and $\sim 0^{0}$ orientation. Table IX shows that the possibility of a conversation satisfying all these constraints is higher near the entrance door compared to the non-entrance door locations. Our volunteer has observed that $\sim$25.3% of 67 people near the entrance satisfied the attack constraints whereas this is 12.3% for non-entrance locations. These are empirical findings and may change with DPS’s locations.
TABLE IX: N = Total people observed, S = Satisfying the constraints in %.
Location Day 1 Day 2 Day 3 Total
#N S% #N S% #N S% #N S%
Near entrance 18 16.6% 28 35.7% 21 19.1% 67 25.3%
Non-entrance 15 6.7% 24 16.6% 18 11.2% 57 12.3%
VIII Countermeasures
Audio damping: The first method is to damp the audio at the input of the pressure ports without affecting the normal pressure measurement. One solution is to use a pressure sampling tube longer than 1 m. Another solution is to enclose the pressure pickup device using a box-like enclosure filled with sound-damping foam to dampen the audio (see Fig. 13 in Appendix E). Both defenses are cheap and easy to adopt.
Reducing sampling frequency of DPSs: Another intuitive solution is to keep the sampling frequency less than 500 Hz. However, reducing sampling frequency is not often possible in some specific applications, where small pressure changes must be detected accurately and quickly for a stable control process. Therefore, for these closely controlled applications, an audio damping device should be used instead of reducing the sampling frequency of pressure sensors.
IX Related Work
Acoustic eavesdropping: Extensive studies on different sensor modalities, such as lasers [46, 47], inertial measurement units (IMU) [48, 49, 50, 51, 13], wireless signals [52, 53, 54, 55, 56, 57], optical sensors [58, 59], vibration motors [60], and hard drives [61] are explored to reveal great threats to speech privacy. mmEcho [52], mmEve [62], and mmSpy [63] use mmWave sensors to capture vibrations and reconstruct speech. However, they cannot reconstruct speech with full intelligibility from the narrowband vibration data. IMU sensors [49, 50, 51] can do digit inference, gender recognition, and limited hot-word reconstruction. However, these works still suffer from the narrowband condition of vibration-based side channels and can only recover band-limited sound with damaged intelligibility. GlowWorm [64] and Lamphone [59] employ optical sensors to recover speech with limited intelligibility. However, they do not recover clean phases, and, as a result, they will not work in transient noisy conditions, like WaLi.
Recently, AccEar [13] and VibSpeech [10] revealed the possibility of using generative adversarial networks to reconstruct intelligible speech with unrestricted vocabulary. WaLi solves the following two problems of AccEar and VibSpeech:
1. Compared to AccEar [13], our WaLi does not require any {audio, pressure sensor data} pair from a specific target victim for training to recover intelligible audio. This makes our attack model more realistic, as it is not practical to have access to the victim’s speech for training.
2. WaLi can recover speech in the presence of transient noise injected from HVAC, whereas, VibSpeech [10] has not been tested for such transient noise. Moreover, WaLi does not require SpkEnc-type encoders and extra vocoders like VibSpeech, as WaLi uses a complex-valued network to extract features from the magnitude and phase of the speech.
Pressure sensor-based eavesdropping: Recently, BaroVox [24] proposed to use DPSs to eavesdrop. The main differences between WaLi and BaroVox are: (1) BaroVox only recognizes hot words or phrases, whereas WaLi can reconstruct intelligible speech with unrestricted vocabulary. (2) BaroVox has not been tested for the transient noise condition. Therefore, it is not realistic for the real-world scenario in HVAC systems.
TABLE X: Comparison between WaLi and recent work.
Comparison Recent works WaLi
Attacking pressure sensors Only [24] does hot word recognition Unrestricted vocabulary
Unrestricted vocabulary [46, 47, 48, 49, 50, 51, 13, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 52, 62, 63] cannot do ✗ Can do ✓
Does not require {audio, pressure data} pair from a target AccEar [13] does not support ✗ Support ✓
Phase reconstruction in noise [10] cannot do ✗ Can do ✓
X Conclusion
We expose a new speech threat that can recover intelligible speech up to 8 kHz from severely aliased pressure sensor data, having a sampling frequency greater than 500 Hz. If a victim unknowingly comes close to the pressure ports and continues a private speech, the attacker can eavesdrop on the victim using our WaLi. WaLi can reconstruct audio even if the speaker is different and is not seen by the model during training time. This makes WaLi more flexible compared to previous work. Using our WaLi, an attacker can secretly listen to natural conversation behind the wall that is the least unexpected. Moreover, we comprehensively evaluate WaLi using six metrics that have not been done before.
Appendix A Pitch, Phonemes, and Intelligible Bandwidth
TABLE XI: Pressure ranges for different voice conditions [15].
Condition in dB in Pa
Quiet whisper 40 dB 0.02 Pa
Conversation 60 dB 0.2 Pa
Loud speech 70-100 dB 0.6-10 Pa
Appendix B Impact of Speech on Pressure Sensors
Table I indicates that DPSs have a sampling frequency of 0.5-2 kHz. However, at least 8 kHz sampling is required to recover intelligible speech. Therefore, a natural question is: Will it be possible to capture intelligible speech with at least 0.5 kHz sampling frequency using pressure sensors?. The pitch frequencies for a male speaker vary between 85-180 Hz, whereas for a female speaker, they vary between 165-255 Hz [65]. Although the 0.5 kHz sampling rate can capture pitches and a few harmonics, it is not sufficient to provide perfect intelligibility [8], as most harmonics at high frequencies will be missing from the bandwidth.
Appendix C Sensor Details
TABLE XII: Sensor details.
Part name Vendor Range
SDP810-125PA Sensiron -125 - +125 Pa
SDP810-500PA Sensiron -500 - +500 Pa
SETRA264 Setra Systems -0.1 inch WC to +0.1 inch WC
Appendix D Details on Comprehensive Evaluation Metrics
We evaluate WaLi using the following six different metrics.
b) Non-Intrusive Speech Quality Assessment - Mean Opinion Score (NISQA-MOS): NISQA-MOS is used to estimate the perceived quality of reconstructed audio without needing a reference signal and gives values between 0 to 5.
e) Short-Time Objective Intelligibility (STOI): STOI predicts how understandable speech is to human listeners in the presence of noise and gives values between 0 to 1.
f) Word Error Rate (WER): WER measures how many words are incorrectly reconstructed compared to a reference transcript [66]. A lower WER means better speech quality.
Figure 13: Long pressure tube and box-like enclosure.
Appendix E Audio damping
The first method (see Fig. 13) is to damp the audio using a damping device at the input of the pressure ports without affecting the normal pressure measurement. One solution is to use a pressure sampling tube longer than 1 m to dampen the audio significantly. Another solution is to enclose the pressure pickup device using a box-like enclosure filled with sound-damping foam to dampen the audio. Both of these countermeasures are cheap and easy to adopt.
Appendix F Comparative Analysis with Recent Work
TABLE XIII: Comparison between WaLi and recent work.
Comparison Recent works WaLi
Attacking pressure sensors Only [24] does hot word recognition Unrestricted vocabulary
Unrestricted vocabulary [46, 47, 48, 49, 50, 51, 13, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 52, 62, 63] cannot do ✗ Can do ✓
Does not require {audio, pressure data} pair from a target AccEar [13] does not support ✗ Support ✓
Phase reconstruction in noise [10] cannot do ✗ Can do ✓
Appendix G Sampling For Less than 500 Hz
@@
Figure 15: The MATLAB interface used in Subjective Tests.
Appendix H Subjective Evaluation Details
To evaluate the performance of the proposed WaLi, a formal listening test has been conducted. A total of 18 undergraduate students with mixed demography who self-reported normal-hearing (NH) participants—comprising ten males and eight females with an average age of 22—participated in the study. The participants voluntarily join to rate the audios without any compensation. At first they are trained on how to assign scores based on perceived perceptual quality of audios. They are also briefed about the purpose of the experiments, potential risks, and about the outcome of this paper. All participants were native English speakers and used soundproof headsets to ensure consistent and distraction-free listening conditions. Each participant evaluated 30 sets of speech samples. Every set included two randomly presented versions of the same utterance: (i) the unprocessed pressure sensor signal, and (ii) the speech processed by the proposed WaLi. Participants rated the perceptual quality of each sample on a 5-point Mean Opinion Score (MOS) scale, where 1 indicates the lowest and 5 the highest quality. The test used speech clips, which were processed under two different frequency band conditions: 500 Hz to 8 KHz and 2 kHz to 8 kHz. In the Fig. 15, we have presented the MATLAB interface that we use to conduct the subjective evaluation.
s